• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 21:40:20 on February 07, 2009

    Entry number 259576


    keyword=Pedestal Suppression File Updated for S2m

    I updated the pedestal suppression file for the S2m detector to reflect
    the ADC swap for the left PMTs. From the online replay, the pedestal
    suppression for the right-side PMTs was too tight. I loosened the
    pedestal suppression for all 16 channels. I have seen for that for these
    16 ADC channels that the pedestals slowly change over time.

    I have kept an eye on it during Transversity. Since mid-November the
    pedestals have drifted lower by about 30 channels. I believe that the
    ADC in ROC 3, Slot 25 is starting to have problems. This is the ADC I
    moved the 16 left S2m channels from yesterday.

    Currently the 16 right S2m PMTs, 10 Cerenkov PMTs and Cerenkov sum are
    still in this ADC. With A1 no longer in use, we could transfer these
    signals into slot 23, which has enough spare channels.

    A copy of this log entry has been emailed to: brads,moffit,rom,camsonne